# Rethinking CLIP-based Video Learners in Cross-Domain Open-Vocabulary Action Recognition

### v1

@@ Abstract
Contrastive Language-Image Pretraining (CLIP) has shown remarkable open-vocabulary abilities across various image understanding tasks. Building upon this impressive success, recent pioneer works have proposed to adapt the powerful CLIP to video data, leading to efficient and effective video learners for open-vocabulary action recognition. Inspired by the fact that humans perform actions in diverse environments, our work delves into an intriguing question: Can CLIP-based video learners effectively generalize to video domains they have not encountered during training? To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, and conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners under various types of domain gaps. Our evaluation demonstrates that previous methods exhibit limited action recognition performance in unseen video domains, revealing potential challenges of the cross-domain open-vocabulary action recognition task. To address this task, our work focuses on a critical challenge, namely scene bias, and we accordingly contribute a novel scene-aware video-text alignment method. Our key idea is to distinguish video representations apart from scene-encoded text representations, aiming to learn scene-agnostic video representations for recognizing actions across domains. Extensive experimental results demonstrate the effectiveness of our method. The benchmark and code will be available at https://github.com/KunyuLin/XOV-Action/.
@@ 1 Introduction
Figure 1: (a) Can CLIP-based video learners effectively generalize to unseen test domains? For example, given a model trained with normal videos, we wonder if it can effectively recognize actions in dark videos. (b) We conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners on four test datasets, namely UCF [59], HMDB [30], ARID [78] and NEC-Dr [12]. For each test dataset, we report the accuracy of closed-set and open-set action categories according to the training categories in Kinetics400 [6]. As shown above, these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Note that, for each metric, we report the best performance among all methods (namely different checkpoints), please refer to Table 3 for the full results. Best viewed in color.
@@
Action recognition aims to recognize what actions humans are performing in videos, which has wide applications in surveillance systems, health monitoring, sport analysis, etc [28, 62, 14, 13, 94, 17, 22]. Recently, inspired by the impressive success of Contrastive Language-Image Pretraining (CLIP) [52] across various image understanding tasks, pioneer works propose to adapt CLIP to video data for action recognition [72, 46, 26, 75, 53, 73]. Different from traditional models that focus on closed-set recognition [56, 71, 38, 6, 3], the new CLIP-based recognition paradigm leads to efficient video learners with remarkable open-vocabulary recognition abilities, i.e., they achieve state-of-the-art performance for both closed-set and open-set action categories on various popular video datasets with moderate training cost. Such open-vocabulary abilities significantly enhance the practical relevance of action recognition models.
@@
As shown in Figure 1(a), in this work, we delve into an intriguing question: Can CLIP-based video learners effectively generalize to video domains that they have not encountered during training? The generalization ability in unseen video domains is crucial for action recognition models in practice, since models often suffer from environment changes when deployed in real-world applications [7, 85]. For example, surveillance systems will encounter actions performed under illumination shifts caused by day-night change or weather change. Therefore, from a practical perspective, we expect that a CLIP-based video learner possessing open-vocabulary abilities can robustly adapt to domain changes for action recognition.
@@
To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, where the four test datasets exhibit various levels of domain gaps in comparison to the training datasets. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners for cross-domain open-vocabulary action recognition. As shown in Figure 1(b), our evaluation reveals that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Specifically, these models exhibit very limited performance for closed-set categories in test domains with large domain gap (e.g., ACC of 53.89% for dark videos in ARID). Also, we note that the open-set recognition performance of these CLIP-based video learners is far from reaching saturation even in domains with moderate domain gap (i.e., ACC of 42.22% on HMDB). Overall, cross-domain open-vocabulary action recognition is challenging and underexplored.
@@
To address this task, our work focuses on tackling the domain gaps by mitigating scene bias, which is a critical challenge of the task. Specifically, the scene bias is caused by the high correlation between human actions and specific scenes in training videos, also known as the spurious correlation [36, 11, 69, 33]. For example, as shown in Figure 1(a), humans usually perform jumping in track and field in the training domain (Kinetics400), thus models prefer to recognize the action according to the track and field since the static scenes are easier to fit [36, 11, 69]. However, humans may perform jumping in hallways (ARID) or basketball courts (NEC-Dr), thus recognizing “jump” by the track and field would result in recognition errors across domains.
@@
To mitigate the scene bias, we propose a novel Scene-Aware video-Text Alignment (SATA) method for learning scene-agnostic video representations. Specifically, we propose to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information and thus pay more attention to action information.
@@
  • We establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, including four test domains exhibiting various levels of domain gap with respect to the training domains. We identify closed-set and open-set categories for each test domain, thus providing a comprehensive way to evaluate models across various situations.
  • We conduct a thorough evaluation of five state-of-the-art CLIP-based video learners based on XOV-Action. Our experimental analysis demonstrates the limited performance of these methods in cross-domain open-vocabulary action recognition, revealing potential challenges of this task.
  • We propose a novel scene-aware video-text alignment method, which learns scene-agnostic video representations to recognize actions across domains. And, extensive experiments demonstrate the effectiveness of our method.
@@ 2.1 Action Recognition
Action recognition aims to recognize human actions in videos. In the past decade, motivated by the success of deep learning [29, 57, 64, 21, 63, 10, 68], many video classification architectures have been proposed. Among these architectures, there are three most popular types, namely 2D CNNs, 3D CNNs and Video Transformers. Typically, 2D CNNs adopt 2D convolution for spatial modeling, and conduct temporal modeling beyond spatial modeling [71, 93] or embed temporal shift into spatial modeling [38, 55, 60]. Different from 2D models, 3D CNNs extends 2D convolution to 3D convolution for adapting video data [65, 6, 66, 67, 16, 90, 35]. By adopting attention mechanisms, Video Transformers expand the receptive field of 2D and 3D CNNs, leading to remarkable performance [19, 45, 89, 91, 3, 2]. Although above models show promising performance for closed-set action recognition, they usually lack the ability to recognize open-set categories.
2.2 Cross-Domain Action Recognition
@@
Cross-domain action recognition aims to learning video classification models by transferring knowledge from source domains to target domains. This area includes two tasks, namely domain-adaptive and domain-generalizable action recognition. In domain-adaptive action recognition [61, 77, 25, 7, 79, 5, 80], unlabeled videos from target domains are accessible for training, and thus related works usually focusing on developing models oriented to specific target domains. Typical works address domain-adaptive action recognition by learning cross-domain invariance, i.e., exploring the shared action cues between the source and target domains [7, 48, 48, 81, 42, 54]. Other works usually focus on leveraging the multi-modal nature of video data [44, 58, 27, 84, 92].
@@
Different from domain-adaptive action recognition, domain-generalizable action recognition [85, 50, 39, 51] aims to learn generalizable models in unseen test domains, i.e., videos of the target domains are not accessible during training. In the absence of target videos, existing works usually assume specific types of invariance across domains for generalizable action representation learning, e.g., Yao et al. assume that local temporal features are more invariant across domains compared with global temporal features [85]. Instead of invariance mining, Lin et al. propose to learn diverse action features for generalization in unseen target domains [39]. Our cross-domain open-vocabulary action recognition task is closely related to domain-generalizable action recognition, since test domains are unseen during training. However, our task is much more challenging as we strive for open-vocabulary abilities, leading to significantly different technical designs.
2.3 Vision-Language Pretraining
@@
Vision-language pretraining [52, 34, 37, 1, 88, 86, 15, 83, 18] has made great progress in recent years, and one of the most impressive works is Contrastive Language-Image Pretraining (CLIP) [52]. By utilizing web-scale paired image-text data for training, CLIP shows robust zero-shot object recognition abilities. Also, some advanced works have demonstrated that CLIP can effectively solve specific downstream tasks by efficient adaptation [97, 87, 49, 96, 53, 41]. Moreover, integrating CLIP with specialized techniques shows remarkable open-vocabulary abilities on various image understanding tasks, such as object detection and segmentation [52, 74, 20, 31, 40]. Although large-scale image-text pretraining has achieved great success, video-text pretraining still has room for development [76, 32, 70, 43, 23, 9, 53]. It is because videos are inherently more complex than images, and large-scale paired video-text datasets are less available. Therefore, it is valuable to develop efficient methods for adapting pretrained image-text models to video understanding tasks.
@@ 3 The Proposed XOV-Action Benchmark
Our work focuses on the cross-domain open-vocabulary action recognition task. This task aims to develop open-vocabulary action recognition models that are generalizable in unseen target domains by training in the source domain, i.e., recognizing both closed-set and open-set action categories in new test domains. This task, closely aligning with real-world applications, requires action recognition models to exhibit open-vocabulary classification abilities and demonstrate robustness in handling environmental changes. To evaluate models on this task, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which provides a comprehensive way to analyze models across various situations. In what follows, we introduce the components of our XOV-Action benchmark in detail, as well as evaluation metrics and potential challenges.
Benchmark Components: Our proposed XOV-Action benchmark consists of two source datasets for training and four target datasets for test. Specifically, the two source datasets used for training are as follows:
(1) Kinetics400 [6]: One of the most widely-used action recognition datasets, consisting of 400 action categories. Videos in Kinetics400 are collected from YouTube, which are usually recorded in normal environments (e.g., normal illumination and weather). Existing works of CLIP-based video learners typically use Kinetics400 for training, and we use the original train-validation split following them.
(2) Kinetics150: A subset of Kinetics400, composed of 150 action categories selected from the full Kinetics400. These 150 categories include all the closed-set categories of the four target datasets in comparison to Kinetics400, and the remaining categories are random sampled. We will illustrate the definition of closed-set categories later. We construct this subset for a detailed analysis to cross-domain open-vocabulary action recognition models.
The four target datasets for test are as follows:
(2) HMDB [30]: A widely-used action recognition datasets consisting of 51 action categories. Compared with UCF, videos in HMDB are captured from more unconstrained environments and more different camera views. Specifically, Videos in HMDB are collected mainly from movies, and remaining videos are from Prelinger archive, YouTube or Google videos. Overall, HMDB has a moderate domain gap compared with the Kinetics domain [80, 82].
(3) ARID [78]: A dataset consisting of 11 categories of action videos, which are recorded under dark environments. These actions include singular person actions (e.g., jump, run) and actions associated with objects (e.g., drink, pick). Due to the significant different illumination conditions, ARID exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example). Developing models that are generalizable in ARID is challenging, when training data are limited to videos with normal illumination.
In our experiments, we use videos from Kinetics400 or Kinetics150 for training. Then, we evaluate models on the four target datasets, which exhibit various levels of domain gap in comparison to the training datasets. Please refer to the Appendix for more details of our XOV-Action.
Evaluation Metrics: First, we illustrate the definitions of closed-set and open-set categories for test domains: (a) The closed-set categories refer to the categories that share similar meanings and have a common lexicon with the categories in the training domain, following a similar approach in previous zero-shot action recognition works [4, 8, 95]. (b) The open-set categories refer to the remaining categories that are not involved in the training domain. Then, we identify closed-set and open-set categories for each test domain, according to the Kinetics400 categories. Category statistics are summarized in Table 1.
In our experiments, we adopt three evaluation metrics:
(1) The closed-set accuracy measures the recognition performance of closed-set categories, which primarily evaluates the model abilities of tackling domain gaps.
(2) The open-set accuracy measures the performance of open-set categories, which evaluates the generalization abilities across both video domains and action categories.
(3) The overall accuracy measures the recognition performance over all categories, which provides a holistic view of model effectiveness across various situations.
Note that, some previous works of CLIP-based video learners treat all the categories in UCF and HMDB as open-set for models trained on Kinetics400. However, UCF and HMDB have many overlapping categories with Kinetics400, thus these works make an inaccurate assessment of the model’s open-vocabulary abilities. Unlike these works, we provide a more accurate way for evaluation by distinguishing between closed-set and open-set categories. By evaluating across various domain gaps and different action categories, our proposed XOV-Action benchmark enables us to perform a wide range of analysis for cross-domain open-vocabulary action recognition models.
Potential Challenges: Although cross-domain open-vocabulary action recognition is valuable and practical, it is challenging and underexplored. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners, as shown in Table 3. The experimental results demonstrate that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Specifically, the results reveal three fundamental challenges of the cross-domain open-vocabulary action recognition task as follows:
(1) How to tackle the domain gap between the training and test domains? Previous CLIP-based video learners exhibit very limited performance in test domains with large domain gaps, i.e., the best closed-set ACCs are only 53.89% and 31.48% on ARID and NEC-Dr, respectively.
(2) How to improve the generalization abilities for open-set categories in the presence of domain gaps? The cross-domain open-set recognition performance is far from reaching saturation even in test domains with moderate domain gaps, e.g., the best open-set ACC is only 42.22% on HMDB.
(3) How to develop a single model that achieves superior performance on all the test domains? It is very challenging since existing models demonstrate varied performance across various test domains as shown in Table 3.
@@
Table 1: Category statistics of four test domains. Closed-set and open-set categories are identified according to Kinetics400.

### v2

@@ Abstract
Building upon the impressive success of CLIP (Contrastive Language-Image Pretraining), recent pioneer works have proposed to adapt the powerful CLIP to video data, leading to efficient and effective video learners for open-vocabulary action recognition. Inspired by that humans perform actions in diverse environments, our work delves into an intriguing question: Can CLIP-based video learners effectively generalize to video domains they have not encountered during training? To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, and conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners under various types of domain gaps. The evaluation demonstrates that previous methods exhibit limited action recognition performance in unseen video domains, revealing potential challenges of the cross-domain open-vocabulary action recognition task. In this paper, we focus on one critical challenge of the task, namely scene bias, and accordingly contribute a novel scene-aware video-text alignment method. Our key idea is to distinguish video representations apart from scene-encoded text representations, aiming to learn scene-agnostic video representations for recognizing actions across domains. Extensive experiments demonstrate the effectiveness of our method. The benchmark and code will be available at https://github.com/KunyuLin/XOV-Action/.
@@ 1 Introduction
Action recognition aims to recognize what actions humans are performing in videos, which has wide applications in surveillance systems, health monitoring, etc [1; 2]. Recently, inspired by the impressive success of CLIP (Contrastive Language-Image Pretraining) [3] across various image understanding tasks, pioneer works propose to adapt CLIP to video data for action recognition [4; 5; 6; 7; 8; 9]. Different from traditional models that focus on closed-set recognition [10; 11; 12; 13; 14], this new CLIP-based recognition paradigm leads to efficient video learners with remarkable open-vocabulary recognition abilities, i.e., they achieve state-of-the-art performance for both closed-set and open-set action categories on various video datasets with moderate training cost. Such open-vocabulary abilities significantly enhance the practical value of action recognition models.
@@
As shown in Figure 1 (a), in this work, we delve into an intriguing question: Can CLIP-based video learners effectively generalize to video domains that they have not encountered during training? This generalization ability in unseen video domains is crucial for action recognition models in practice, since models often suffer from environment changes when deployed in real-world applications [15; 16]. For example, surveillance systems will encounter actions performed under illumination shifts caused by day-night change or weather change. Therefore, from a practical perspective, we expect that a CLIP-based video learner possessing open-vocabulary abilities can robustly adapt to domain changes for action recognition.
Figure 1: (a) Can CLIP-based video learners effectively generalize to unseen test domains? For example, given a model trained with normal videos, we wonder if it can effectively recognize actions in dark videos. (b) We conduct a comprehensive evaluation for five state-of-the-art CLIP-based video learners on four test datasets, namely UCF [17], HMDB [18], ARID [19] and NEC-Dr [20]. For each test dataset, we report the accuracy of closed-set and open-set action categories according to the training categories in Kinetics400 [13]. As shown above, these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Note that, for each metric, we report the best performance among all methods, please refer to Table 3 for the full results.
@@
To answer this question, we establish the first CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which includes four test datasets exhibiting various levels of domain gaps in comparison to the training datasets. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners in cross-domain open-vocabulary action recognition. As shown in Figure 1 (b), our evaluation reveals that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Specifically, these models exhibit very limited performance for closed-set categories in test domains with large domain gap (e.g., ACC of 53.89% for dark videos in ARID). Also, we note that the open-set recognition performance of these CLIP-based video learners is far from reaching saturation even in domains with moderate domain gap (i.e., ACC of 42.22% on HMDB). Overall, cross-domain open-vocabulary action recognition poses challenges for existing CLIP-based video learners, as it requires recognizing both closed-set and open-set action categories in unseen video domains.
@@
To address this task, our work focuses on tackling domain gaps by defending against the scene bias when fitting training videos. Specifically, the scene bias is caused by the strong associations between actions and specific scenes in training videos (aka., spurious correlation [21; 22; 23; 24; 25; 26]), and it would hinder models’ generalizability across domains. For example, as shown in Figure 1 (a), if humans usually perform jumping on track-and-field grounds in the training domain (Kinetics400), models will prefer to recognize the action based on the track-and-field grounds since static scenes are easier to fit [21; 22; 24]. However, humans may perform jumping in hallways (ARID) or on basketball courts (NEC-Dr), thus recognizing “jump” based on track-and-field grounds would result in recognition errors across domains.
@@
To mitigate the scene bias, we propose a novel Scene-Aware video-Text Alignment (SATA) method for learning scene-agnostic video representations. Our key idea is to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information and thus pay more attention to action information. Experiments on XOV-Action demonstrate that our SATA can effectively improve the closed-set action recognition performance across domains while generally maintaining the open-set performance.
@@
(1) We establish the first CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which includes four test domains exhibiting various levels of domain gap in comparison to the training domains. We identify closed-set and open-set categories for each test domain, thus providing a comprehensive way to evaluate models across various situations.
(2) We conduct a thorough evaluation of five state-of-the-art CLIP-based video learners based on XOV-Action. Our experimental analysis demonstrates the limited performance of these methods in cross-domain open-vocabulary action recognition, revealing potential challenges of this task.
(3) We propose a novel Scene-Aware video-Text Alignment (SATA) method, which aims to learn scene-agnostic video representations to recognize actions across domains. Extensive experiments demonstrate the effectiveness of our method in tackling domain gaps.
@@ 2 Related Works
Action Recognition aims to recognize human actions in videos. In the past decade, motivated by the success of deep learning [27; 28; 29; 30; 31; 32; 33], many video classification architectures have been proposed. These architectures can primarily be categorized into 2D CNNs, 3D CNNs and Video Transformers. Typically, 2D CNNs adopt 2D convolution for spatial modeling, and conduct temporal modeling beyond spatial modeling [11; 34] or embed temporal shift into spatial modeling [12; 35; 36]. 3D CNNs extends 2D convolution to 3D convolution for adapting video data [37; 13; 38; 39; 40; 41; 42]. By adopting attention mechanisms, Video Transformers expand the receptive field of 2D and 3D CNNs, leading to remarkable performance [43; 44; 45; 46; 14; 47]. Although above models show promising performance for closed-set action recognition, they usually lack the ability to recognize open-set categories.
Recently, inspired by the success of CLIP (Contrastive Language-Image Pretraining) [3; 48], some pioneer works propose to adapt CLIP to video data for action recognition [49; 4; 5; 49; 6; 7; 8; 9; 50; 51; 52; 53; 54; 55]. Owing to the image-text alignment power of CLIP, these CLIP-based video learners are endowed with remarkable open-vocabulary action recognition abilities with moderate training cost. These methods adapt CLIP to video data in various ways. For example, ActionCLIP [4] stacks temporal fusion layers on top of the image encoder for modeling temporal dynamics, Ju et al. [6] cooperate continuous prompting with temporal Transformer, and X-CLIP [5] proposes cross-frame communication attention for temporal modeling. OpenVCLIP [9] and FROSTER [54] propose to improve open-set recognition by harnessing the power of raw CLIP. In our work, we focus on the underexplored cross-domain setting for open-vocabulary action recognition, which aims to develop generalizable open-vocabulary action recognition models for unseen video domains.
@@
Cross-Domain Action Recognition aims to learning video classification models by transferring knowledge from source domains to target domains. This area mainly includes two tasks, namely domain-adaptive and domain-generalizable action recognition. In domain-adaptive action recognition [56; 57; 58; 15; 59; 60; 61], unlabeled videos from target domains are accessible for training, thus prevailing works usually focusing on developing models oriented to specific target domains. Typical works address domain-adaptive action recognition by learning cross-domain invariance [15; 62; 62; 63; 64; 65]. Other works usually explore to leverage the multi-modal nature of video data [66; 67; 68; 69; 70].
@@
Differently, domain-generalizable action recognition [16; 71; 72; 73] aims to learn generalizable models in unseen test domains, i.e., videos of the target domains are not accessible during training. In the absence of target videos, existing works usually assume specific types of invariance across domains for generalizable action representation learning, e.g., Yao et al. assume that local features are more invariant across domains compared with global features [16]. Our cross-domain open-vocabulary action recognition task is closely related to domain-generalizable action recognition, since test domains are unseen during training. However, our task is much more challenging as we strive for cross-domain open-vocabulary abilities, leading to significantly different technical designs.
@@
Vision-Language Pretraining [3; 48; 74; 75; 76; 77; 78] has made great progress in recent years, and one of the most impressive works is CLIP [3]. By utilizing web-scale paired image-text data for training, CLIP shows robust zero-shot object recognition abilities. Also, some advanced works have demonstrated that CLIP can effectively solve specific downstream tasks by efficient adaptation [79; 80; 81; 82; 8; 83]. Moreover, integrating CLIP with specialized techniques shows remarkable open-vocabulary abilities on various image understanding tasks, e.g., object detection and segmentation [84; 85; 86; 49]. Although large-scale image-text pretraining has achieved great success, video-text pretraining still has room for development [87; 88; 89; 90; 91; 92; 93; 94; 95]. It is because videos are inherently more complex than images, and large-scale paired video-text datasets are less available. Therefore, it is valuable to develop methods for adapting pretrained image-text models to video understanding tasks.
@@ 3 The Proposed XOV-Action Benchmark
Our work focuses on the cross-domain open-vocabulary action recognition task. This task aims to develop open-vocabulary action recognition models that are generalizable in unseen target domains by training in the source domain, i.e., recognizing both closed-set and open-set action categories in new test domains. To evaluate models in this task, we establish XOV-Action, a CROSS-domain Open-Vocabulary Action recognition benchmark, which provides a comprehensive way to analyze models across various situations. In what follows, we introduce the components of our XOV-Action benchmark in detail, as well as evaluation metrics and potential challenges.
@@
Table 1: Category statistics of four test domains in our XOV-Action. The closed-set and open-set categories are identified according to Kinetics400.
@@
Benchmark Components: Our proposed XOV-Action benchmark consists of two source datasets for training and four target datasets for test. The two source datasets for training are as follows:
(1) Kinetics400 [13]: One of the most widely-used action recognition datasets, consisting of 400 action categories. Videos in Kinetics400 are collected from YouTube, which are usually recorded in normal environments (e.g., normal illumination and weather). Existing CLIP-based video learners typically use Kinetics400 for training, and we use the original split following them.
(2) Kinetics150: A subset of Kinetics400, composed of 150 action categories selected from the full Kinetics400. These 150 categories include all the closed-set categories of the four target datasets in comparison to Kinetics400, and the remaining categories are randomly sampled. We will illustrate the definition of closed-set categories later. We construct this subset to conduct detailed analysis for cross-domain open-vocabulary action recognition models.
The four target datasets for test are as follows:
(2) HMDB [18]: A widely-used action recognition datasets consisting of 51 action categories. Compared with UCF, videos in HMDB are captured from more unconstrained environments and more different camera views. Specifically, Videos in HMDB are collected mainly from movies, and remaining videos are from Prelinger archive, YouTube or Google videos. Overall, HMDB has a moderate domain gap compared with the Kinetics domain [61; 96].
(3) ARID [19]: A dataset consisting of 11 categories of action videos, which are recorded under dark environments. These actions include singular person actions (e.g., jump, run) and actions associated with objects (e.g., drink, pick). Due to the significantly different illumination conditions, ARID exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example). When training data are limited to videos with normal illumination, developing models that are generalizable in ARID is challenging.
In our experiments, we use videos from Kinetics400 or Kinetics150 for training. Then, we evaluate models on the four target datasets, which exhibit various levels of domain gap in comparison to the training datasets. Please refer to the Appendix for more details of our XOV-Action benchmark.
Evaluation Metrics: First, we illustrate the definitions of closed-set and open-set categories for test domains: (a) The closed-set categories refer to the categories that share similar meanings and have a common lexicon with the categories in the training domain, following a similar approach in previous zero-shot action recognition works [97; 98]. (b) The open-set categories refer to the remaining categories that are not involved in the training domain. Then, we identify closed-set and open-set categories for each test domain, according to the Kinetics400 categories. Category statistics are summarized in Table 1.
In our experiments, we adopt three evaluation metrics: (1) The closed-set accuracy measures the recognition performance of closed-set categories, which primarily evaluates the model abilities of tackling domain gaps when fitting training videos. (2) The open-set accuracy measures the performance of open-set categories, which evaluates the generalization abilities across both video domains and action categories. (3) The overall accuracy measures the recognition performance over all categories, which provides a holistic view of model effectiveness across various situations.
Note that, some previous works of CLIP-based video learners treat all the categories in UCF and HMDB as open-set for models trained on Kinetics400. However, UCF and HMDB have many overlapping categories with Kinetics400, thus these works make an inaccurate assessment of models’ open-vocabulary abilities. Unlike these works, we provide a more accurate way for evaluation by distinguishing between closed-set and open-set categories. By evaluating across various domains and different action categories, our proposed XOV-Action benchmark enables us to perform a wide range of analysis for cross-domain open-vocabulary action recognition models.
Potential Challenges: Although cross-domain open-vocabulary action recognition is valuable and practical, it is challenging and underexplored. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners. As shown in Table 3, the experimental results on XOV-Action demonstrate that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. In addition, the results reveal three fundamental challenges of the cross-domain open-vocabulary action recognition task as follows:
I. How to improve the generalization of closed-set categories in unseen test domains? Previous CLIP-based video learners exhibit very limited performance in test domains with large domain gaps, i.e., the best closed-set ACCs are only 53.89% and 31.48% on ARID and NEC-Dr, respectively. To address this challenge, it is crucial to learn domain-invariant video representations from training domains. Our method mainly targets this challenge by focusing on the mitigation of scene bias.
II. How to improve the generalization of open-set categories in the presence of domain gaps? The cross-domain open-set recognition performance is far from reaching saturation even in test domains with moderate domain gaps, e.g., the best open-set ACC is only 42.22% on HMDB. To address this challenge, it is crucial to harness the power of language models for recognizing open-set categories across domains. Additionally, another possible direction is to alleviate overfitting of closed-set categories in training domains [9; 54].
III. How to develop a single model that performs excellently for all categories in all test domains? This is very challenging, since existing models show varied performance across various domains, and empirically improvements in cross-domain closed-set performance do not guarantee improvements for open-set categories and could even cause a decrease. Given the unique characteristics of each test domain and category type, it is crucial to find a trade-off solution, especially when no additional annotations/data are introduced and the model size remains consistent.